OpenStudy (anonymous):

What are the factors of x^2 - 5 + 6 a. (x-6) and (x + 3) b. (x + 2) and (x + 3) c. (x + 6) and (x - 3) d. (x - 2) and (x - 3)

OpenStudy (kc_kennylau):

Try to multiply each of the options together :)

OpenStudy (kc_kennylau):

Or if you can, try to name two numbers that times up to \(6\) and adds up to \(-5\) :)

OpenStudy (jojo4eva):

what you want to do is find somethin that when added together equal -5 and when multiplied equal positive 6

OpenStudy (anonymous):

2 * 3 = 6 & -10 + 5

OpenStudy (jojo4eva):

Therefore 2 negatives equal a positive so D is your final answer

OpenStudy (kc_kennylau):

they must be the same number...

OpenStudy (kc_kennylau):

like \(xy=6\) and \(x+y=-5\)

OpenStudy (jojo4eva):

-3-2=-5 and -3*-2=6

OpenStudy (jojo4eva):

So your final answer is D

OpenStudy (skullpatrol):

$$\Huge x^2 - 5 + 6 =x^2+1$$

OpenStudy (anonymous):

?

OpenStudy (kc_kennylau):

?

OpenStudy (anonymous):

I'm confused.

OpenStudy (kc_kennylau):

me too highfive @Yolo4mecuite

OpenStudy (skullpatrol):

Check your question @Yolo4mecuite

OpenStudy (anonymous):

Ok

OpenStudy (kc_kennylau):

oh lol

OpenStudy (anonymous):

I still don't see what he did?

OpenStudy (skullpatrol):

What are the factors of x^2 - 5 + 6

OpenStudy (anonymous):

Oh

OpenStudy (skullpatrol):

You missed an x

OpenStudy (kc_kennylau):

@Yolo4mecuite he's being sarcastic to point out your typo lol

OpenStudy (anonymous):

Oh yeah I am missing an X

OpenStudy (anonymous):

x^2 - 5x + 6

OpenStudy (skullpatrol):

Not sarcastic, you have to learn to $$\Huge \text{read the problem carefully}$$

OpenStudy (anonymous):

Sorry

OpenStudy (kc_kennylau):

@skullpatrol \(\Large\color{red}{\mbox{LOL OK}}\) xD

OpenStudy (anonymous):

so how would I answer the question

OpenStudy (kc_kennylau):

As I said, find out two numbers that multiplies to 6 AND adds up to -5 (or add up to 5 and then make the two numbers negative so that they add up to -5)

OpenStudy (anonymous):

its -2 and -3

OpenStudy (kc_kennylau):

exactly

OpenStudy (anonymous):

Ohh thanks :)

OpenStudy (kc_kennylau):

no problem :)